OpenStudy (anonymous):

Solve the system of equations and enter the solution as an ordered pair. 4x - 4y + 7 = 15 8y = 2x + 14

OpenStudy (amistre64):

is that right? no typos?

OpenStudy (anonymous):

the heat is on....

OpenStudy (amistre64):

4 -4 = 8 -2 8 = 14 according to the cramer rule, as if names were important we can do this: y = (56+16)/(32+8) = 72/40 = 9/5 x = (-56-64)/(8-32) = (-7-8)/(1-4) = -15/3 = -5 if i did it right

OpenStudy (amistre64):

(56+16)/(32+8) /8 /8 ---------------- 7+2 / 4+1 9/5 just chking something out with another personality lol

OpenStudy (amistre64):

8(9)/5 = 2(-5) + 14 72/5 = -10 + 14 ...... it looks peculiar whaich makes me think i forgot how to add someplace

OpenStudy (amistre64):

4x -4y = 8 -2x+ 8y = 14 y = 56 + 16/ 32+8 y = 72/40 y = 9/5 -4y +4x = 8 8y -2x = 14 x = -56-64 / 8-32 x = -120 / -24 = 30 / 6 = 5

OpenStudy (anonymous):

8y=2x+14 Divide by 2, and get: 4y=x+7 4x-4y+7=15 Plug in (x+7) instead of 4y below: 4x-(x+7)+7=15 3x=15 x=5 4y=x+7 4y=5+7=12 y=3

OpenStudy (anonymous):

x=5 y=3

OpenStudy (amistre64):

yeah, but what am i missing ... :) cant practice my cramer if i keep missing it

OpenStudy (anonymous):

cramer vs cramer ?

OpenStudy (amistre64):

no soup for you!

OpenStudy (anonymous):

I don't remember this cramer rule - looking it up now :)

OpenStudy (amistre64):

1x -1y = 2 -1x+ 4y = 7 y = 7(1)- -1(2)/ 4(1)- -1(-1) = 7 + 2 / 4 - 1 = 9/3 = 3 yep, i forgot how to add

OpenStudy (amistre64):

its the determinate of the matrix dividing the alomst determinate of the other thing

OpenStudy (amistre64):

x y ---- a b = m c d = n y = na-cm / ad-bc x = nb - md / bc-ad

OpenStudy (anonymous):

that's almost like cheating :) yeah, it's starting to come back to me now.

OpenStudy (amistre64):

almost; as long as you can remember how to add lol